Strama, Inc., manufactures and sells two products: Product A6 and Product I5. The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product A6 Product I5 Total
Labor-related DLHs $ 189,138 8,000 2,600 10,600
Production orders orders 43,624 800 1,000 1,800
Order size MHs 181,147 5,400 5,500 10,900
$ 413,909

The activity rate for the Order Size activity cost pool under activity-based costing is closest to:

rev: 03_25_2018_QC_CS-119201

Multiple Choice

  • $21.35 per MH

  • $72.75 per MH

  • $39.05 per MH

  • $16.62 per MH

Merone Company allocates materials handling cost to the company's two products using the below data:

Modular Homes Prefab Barns
Total expected units produced 5,200 8,200
Total expected material moves 520 120
Expected direct labor-hours per unit 720 220

The total materials handling cost for the year is expected to be $221,920.

If the materials handling cost is allocated on the basis of direct labor-hours, the total materials handling cost allocated to the prefab barns is closest to: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

  • 121,047.27

  • $56,295.88

  • $123,475.00

  • $72,160.00

Forner, Inc., manufactures and sells two products: Product Z1 and Product Z8. The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product Z1 Product Z8 Total
Labor-related DLHs $ 127,800 4,800 2,400 7,200
Machine setups setups 60,270 1,300 200 1,500
Order size MHs 942,670 3,900 3,500 7,400
$ 1,130,740

The activity rate for the Order Size activity cost pool under activity-based costing is closest to:

rev: 03_25_2018_QC_CS-119201

Multiple Choice

  • $60.25 per MH

  • $185.32 per MH

  • $135.53 per MH

  • $127.39 per MH
Answer #1

1) Activity rate for the order size activity cost pool

Order size = Estimated overhead/estimated machine hour = 181147/10900 = 16.62

So answer is d) $16.62 per MH

2) Estimated labor hour = 5200*720+8200*220 = 5548000

Material handling cost allocated to Pref Barns = 221920/5548000*1804000 = 72160

So answer is d) $72160

3) Activity rate for the order size activity cost pool

Order size = Estimated overhead/estimated machine hour = 942670/7400 = 127.39 per MH

So answer is d) $127.39 per MH